OpenStudy (anonymous):

integrate sin^3x cos^5x dx

OpenStudy (anonymous):

\[\int \sin^3(x)\cos^5(x)dx=-\frac{\cos^6(x)\sin^2(x)}{8}-\frac{\cos^6(x)}{24}\]

myininaya (myininaya):

i cant believe you are here satellite it is early

myininaya (myininaya):

yes looks good :)

myininaya (myininaya):

damn wrong file

myininaya (myininaya):

that is the right one sorry

OpenStudy (anonymous):

which one its right?

myininaya (myininaya):

i know my is right i haven't checked satelliltes

OpenStudy (anonymous):

so qs. when i started doing it i broke up the cos^5x because i have understood that odd exponents u break and use substitution.... what is the exception in this excersice

myininaya (myininaya):

? do you not understand what i did

OpenStudy (anonymous):

no

myininaya (myininaya):

i broke up sin^2x

OpenStudy (anonymous):

u broke uo sin^3x

myininaya (myininaya):

so you don't know sin^3x=sin^2x*sinx

OpenStudy (anonymous):

and rewrite it as cosx

myininaya (myininaya):

right

OpenStudy (anonymous):

yea thats perfect

OpenStudy (anonymous):

i understood that part

OpenStudy (anonymous):

i am talking about cos^5x

OpenStudy (anonymous):

i thought u do the samething

myininaya (myininaya):

you can try it that way if you want

OpenStudy (anonymous):

oh ok.. should give me the same thing?

myininaya (myininaya):

yes if it works

myininaya (myininaya):

yes that will work to i will show you both ways k?

OpenStudy (anonymous):

yea,,, i finished mine.. :) it gave me the same

OpenStudy (anonymous):

i started doing it but i got confuse... maybe i was wrong...

myininaya (myininaya):

why did you get confuse

myininaya (myininaya):

because it was long?

myininaya (myininaya):

by the way satellite is wrong i check his and did not get sin^3xcos^5x back when i took derivative of his anti derivative

OpenStudy (anonymous):

yea mine was longer...lol complicated....

myininaya (myininaya):

oh wait he might be right too i took my algebra alittle bit further if i didn't make a mistake his answer is good

OpenStudy (anonymous):

so how can u get two different answers?

myininaya (myininaya):

it is the same answer

myininaya (myininaya):

just like if we simplifed 6/2 to 3

myininaya (myininaya):

do you remember trig identities

OpenStudy (anonymous):

yes

myininaya (myininaya):

there is alot of ways to right one thing

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

well because i did not see the step that he follow i prefer stick with this ones,,, later when i fell more confortable with this way i can try other... maybe is easier

myininaya (myininaya):

he might had used a formula

OpenStudy (anonymous):

ohhhhhhhhh jijij,,

myininaya (myininaya):

i don't know of the formula or if there is one i could look up and see

OpenStudy (anonymous):

noo,,, i like this way for now.. thanks u myininaya.. very helpful

myininaya (myininaya):

i dont like formulas

myininaya (myininaya):

too easy to use

OpenStudy (anonymous):

i dont like neither

OpenStudy (anonymous):

i am good with numbers,,,, (logic) no with memorization

myininaya (myininaya):

by the way i see those rules you are mentioning now

myininaya (myininaya):

since both were odd it didn't matter what substituition we chose (u=sinx or u=cosx) if the odd was on the cosine though and not the sine, then use u=sinx if the odd was on the sine and not the cosine, then use u=cosx if both powers were even, then you will have to use some trig identities

myininaya (myininaya):

yep yep there is a formula

OpenStudy (anonymous):

oh jijijii

myininaya (myininaya):

this formula would be ugly for really high powers

OpenStudy (anonymous):

u found it?

myininaya (myininaya):

yep

myininaya (myininaya):

i was trying to find a link just in case you wanted to look at it because i didn't want to type it i am still looking online for it, but i have in my cal book in the back

myininaya (myininaya):

\[\int\limits_{}^{}\sin^n(u)\cos^m(u)du=\frac{\sin^{n-1}(u)\cos^{m+1}(u)}{n+m}+\frac{n-1}{n+m}\int\limits_{}^{}\sin^{n-2}(u)\cos^{m}udu\]

OpenStudy (anonymous):

In a general case to integrate stuff of the form sin^nx cos^mx , 1- if both m and n are odd, substitute either sin x=t or cos =t and then integrate. 2- if one is odd and the other is even, sustitute the even power, for eg if m is odd put sin x=t 3- if both are even, dont substitute you can solve by trignometric identities.